OpenStudy (anonymous):

Please help i will give medal and fan. Solve: (-5) x (+4) x (-3) x (+2)

Directrix (directrix):

Are the "x" symbols for multiplication or for the variable x?

OpenStudy (anonymous):

x=multiplication

Directrix (directrix):

Oh, okay.

OpenStudy (anonymous):

please help

Directrix (directrix):

(-5) * (+4) * (-3) * (+2) =

Directrix (directrix):

What is -5 times 4 ? @American.Eagle

OpenStudy (anonymous):

34

OpenStudy (anonymous):

oh sorry i mean -20

Directrix (directrix):

No. 5 times 4 = 20 and then there is the negative, so -20

OpenStudy (anonymous):

okay

Directrix (directrix):

What is this: (-3) * (+2) negative 3 times 2 = ?

OpenStudy (anonymous):

-6

Directrix (directrix):

Right. And finally, what is -6 times -20 ?

Directrix (directrix):

@American.Eagle

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

120

Directrix (directrix):

Okay, What is 6 times 20

OpenStudy (anonymous):

120

Directrix (directrix):

And, a negative times a negative is positive, so the final answer is 120.

OpenStudy (anonymous):

is that the answer

Directrix (directrix):

Do you agree?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think it should be 360

OpenStudy (anonymous):

cuz

Directrix (directrix):

Why?

OpenStudy (anonymous):

-5 * -3 =15 4 * 2 = 8 15 * 8 = 120

OpenStudy (anonymous):

and what was ur ans

OpenStudy (anonymous):

oh i am very sorry u also told 120

OpenStudy (anonymous):

and thank u very much and for ur respect i will give u medal

OpenStudy (anonymous):

can i friend u ?

OpenStudy (anonymous):

i mean fan

Directrix (directrix):

I did it by parts. (-5) * (+4) = -20 (-3) * (+2) = -6 -20 * -6 = + 120 or 120

OpenStudy (anonymous):

okay

Directrix (directrix):

Of course.

Directrix (directrix):

I am already your fan. Now, you will be mine. Cool.

OpenStudy (anonymous):

thanks u very much now i would not fail in my test

OpenStudy (anonymous):

how many medals do u have ?

Directrix (directrix):

Alrighty, then. Every learner and every helper should be respectful, I think. So, we agree on that.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how many medals u have

Directrix (directrix):

You can read that information here: http://openstudy.com/users/directrix

OpenStudy (anonymous):

pellet how u got so many medals. u really are very helpful

OpenStudy (anonymous):

i have one more ques. can i ask

Directrix (directrix):

If it is math, yes.

OpenStudy (anonymous):

(+18) / (+6) * (-3) * (+2)

OpenStudy (anonymous):

/ = division

Directrix (directrix):

Okay, let me think.

Directrix (directrix):

Do you want to say the answer you cranked out now?

OpenStudy (anonymous):

i couldnt get u

Directrix (directrix):

(+18) / (+6) = ? What is 18 divided by 6 = ?

OpenStudy (anonymous):

3

Directrix (directrix):

Okay. What is (-3) * (+2) ? What is 3 times 2 and then add the negative symbol?

OpenStudy (anonymous):

-6

Directrix (directrix):

And, last what is 3 * (-6)? Three times six and add on the negative symbol.

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

thanks now i am done

Directrix (directrix):

Correct, that is what I got.

OpenStudy (anonymous):

how to know that if somene is online or not

Directrix (directrix):

If you have other questions, close this thread and post the other questions in a new thread.

Directrix (directrix):

Hover your cursor over the name of the person. You will see something like this attachment.

Directrix (directrix):

I don't know of any direct way to search for somebody online.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

and keep it up if u get one more medal u would totally get 6000 medals

Directrix (directrix):

I will try. Thank you.

OpenStudy (anonymous):

gtg

OpenStudy (anonymous):

bye

Directrix (directrix):

See you when you are on OS later.